OpenStudy (anonymous):

What is the solution set of the given equation? 4y - 6y + 9y = -2 {-2/11} {-2/7} {2/7} {2/11}

OpenStudy (shamil98):

Combine the like terms of y and then divide both sides to isolate y.

OpenStudy (shamil98):

4 - 6 + 9 = what?

OpenStudy (anonymous):

11?

OpenStudy (shamil98):

-2 + 9 remember pemdas you solve from left to right :)

OpenStudy (shamil98):

7y = -2

OpenStudy (shamil98):

Now you want to divide both sides by 7 to isolate y and get your answer.

OpenStudy (anonymous):

huh?

OpenStudy (shamil98):

4y - 6y + 9y = -2 7y = -2 y = -2/7..

OpenStudy (anonymous):

Solution is. -2/7

OpenStudy (anonymous):

thanx guys